OpenStudy (anonymous):

3logx=2 I need someone to explain what Im suppose to do??

OpenStudy (anonymous):

solve for x. 3*log x = 2. divide by 3 on both sides log x = 2/3. solve for x.

OpenStudy (anonymous):

the value of x depends on your base.

OpenStudy (anonymous):

What do I need to do to solve for x?

OpenStudy (anonymous):

\[\log_{e}x = y \rightarrow x = e ^{y}\]

OpenStudy (anonymous):

so there is no answer??

OpenStudy (anonymous):

of course there is. I explained how to get x from log x in the previous post.

OpenStudy (anonymous):

Thanks for your help. I still dont understand, its just me. Thank you.

OpenStudy (anonymous):

log x = 2/3 ==> x = base^2/3. if it is natural log, then base is e. But in this case i think base is 10. so x = 10^2/3

OpenStudy (anonymous):

Is there something I should be punching in my caculator to get this answer?

OpenStudy (anonymous):

type 10^(2/3) in google

OpenStudy (anonymous):

copy and paste that exact phrase

OpenStudy (anonymous):

I did that and its giving me an answer of 4.??. I just wanted to know from logx=2/3 if there was something I needed to be doing on my cac. Thanks again.

OpenStudy (anonymous):

I dont know what calculator you have. Refer to the calculator instructions.

OpenStudy (anonymous):

I understand that. I was just looking for some basic steps to get from logx=2/3 to the answer of 10^2/3. Because I dont know what to do from there. I dont even know what to refer to in my cac, because Im not sure of the next step. BUR, I feel really really dumb, and Im sure Im just bothering you. Thanks